OpenStudy (anonymous):

a great proof of 1=2

OpenStudy (anonymous):

Let ... a = b Multiply both sides by 'a' Add a2 to both sides Subtract 2ab from both sides cancel the common factor from both sides Hence 2=1

OpenStudy (anonymous):

Think and tell what is done wrong here ?

OpenStudy (anonymous):

Let a=b multiply both sides by 'a' a2=ab add a2 to both sides a2+a2=a2+ab 2a2=a2+ab Subtract 2ab from both sides 2a2-2ab=a2+ab-2ab 2a2-2ab=a2-ab 2(a2-ab)=1(a2-ab) therefore 1=2...........

OpenStudy (anonymous):

Let a = b multiply both sides by a a^2 = a*b subtract b^2 from both sides a^2-b^2 = a*b-b^2 apply the distributive law to both sides (a+b)(a-b) = b(a-b) divide both sides by (a-b) (a+b) = b substitute all a's for b's (remember, if a = b you can do this)/ a+a = a regroup the two a's in the left side, and rename it 2a 2a = a divide both sides by a 2 = 1

OpenStudy (anonymous):

2(a2-ab)=1(a2-ab) how can u cancel while a2-ab=0

OpenStudy (anonymous):

0/0 is undefined

OpenStudy (anonymous):

wrong step

OpenStudy (anonymous):

any more answers !

OpenStudy (anonymous):

you can not devide a number by zero. :))

OpenStudy (anonymous):

ok any more ?

OpenStudy (anonymous):

Sajan And zakaullah is right

OpenStudy (kira_yamato):

Let f(x) = x+1 and g(x) = x f(1) = 2, g(1) = 1 f'(x) = 1, g'(x) = 1 f'(x) = g'(x) f(x) = g(x) f(1) = g(1) 2 = 1